# Rethinking Attention Output Projection: Structured Hadamard Transforms for Efficient Transformers

### 09 Mar 2026 v1

@@ Abstract
The dense output projection in multi-head attention scales quadratically with model dimension, contributing significantly to parameter count, memory footprint, and inference cost. We propose replacing this projection with a fixed, parameter-free Walsh–Hadamard Transform followed by a lightweight learnable affine rescaling, eliminating approximately 25% of attention parameters per block while preserving global cross-head interaction through an orthogonal, norm-preserving transformation. Across different model sizes, we demonstrate that this structured substitution maintains comparable or slightly superior downstream task performance on standard benchmarks, while achieving up to 7% aggregate parameter reduction, 8.9% peak memory savings, and 6.6% throughput improvement at scale, with efficiency gains growing monotonically with model size, batch size, and sequence length. Interestingly, we observe that structured Hadamard-based models exhibit a steeper validation loss curve relative to training FLOPs compared to their dense counterparts, suggesting more favorable compute utilization during training.
@@ 3 Motivation
The Hadamard matrix presents a principled answer to this question. As a fixed, parameter-free orthogonal transform, the Walsh–Hadamard transform (WHT) uniformly mixes all input dimensions through a butterfly-structured computation (Figure 2), requiring no learned weights while preserving the norm of its input. Crucially, this structured mixing is not merely a computational convenience—it imposes an inductive bias on how heads interact. Because the Hadamard transform couples every head output with every other through a fixed, maximally spread orthogonal basis, it encourages the model to distribute information across heads in a globally coherent manner. Rather than permitting arbitrary redundancy through unconstrained linear recombination, this structure implicitly regularizes the attention module: heads are incentivized to learn complementary, non-overlapping representations, since only such representations can be efficiently preserved and distinguished under a fixed orthogonal mixing. This is analogous in spirit to the role of orthogonal initialization in deep networks structured constraints that promote representational diversity without explicit supervision.
From a computational perspective, the WHT admits an butterfly factorization, replacing the dense matrix multiply with a cascade of lightweight addition and subtraction operations across stages. Replacing the output projection with this transform eliminates a parameter matrix that typically accounts for roughly one-quarter of the total attention parameters in a standard MHA block, without introducing any new hyperparameters or architectural complexity.
Together, these properties—parameter efficiency, orthogonality, inductive bias toward complementary head representations, and computational regularity—motivate the use of the Hadamard transform as a structured drop-in replacement for the dense output projection in multi-head attention.
In this work, we propose to replace the dense output projection in attention block with a structured transformation based Hadamard matrix. By doing so, we preserve the essential role of head interaction and dimensional consistency required for residual connections, while significantly reducing the number of learnable parameters. The resulting formulation removes a dense parameter matrix whose size scales quadratically with the model dimension, leading to a reduction of approximately one quarter of the total attention parameters in a standard multi-head attention block. In figure 2, we have shown the flow of the fast Walsh-Hadamard transform butterfly diagram, illustrating the three-stage computation for n=8 inputs. This structured approach challenges the assumption that full dense mixing is necessary for effective head aggregation, suggesting that a fully dense projection may be unnecessarily overparameterized and offering a more parameter-efficient alternative.
@@ 4.1 Architecture Overview
We propose a modified multi-head attention (MHA) block in which the dense output projection is replaced by a structured Hadamard-based mixing operation Dao-AILab Contributors (2022). All other components—query, key, value projections, and per-head attention computation—remain unchanged.
Given input , standard MHA computes:
and produces concatenated head outputs , which are then mixed via a learned dense projection .
We replace this projection entirely with a fixed orthogonal Hadamard transform , satisfying , followed by a learnable affine rescaling:
where are trainable scale and bias parameters. (H is orthogonal i.e 1/n-embed)
The motivation is that full dense coupling across attention heads, while expressive, introduces quadratic parameter overhead and potential redundancy. The Hadamard transform achieves global, information-preserving cross-head interaction through additions and sign changes alone, substantially reducing parameter count while retaining sufficient expressive capacity through the affine parameters.
@@ 4.2 Parameter Efficiency Analysis
The attention block in standard MHA contains four dominant parameter matrices: query, key, value, and output projections, each contributing parameters. Ignoring bias terms, the total number of attention parameters is therefore
with the output projection alone accounting for parameters.
In the proposed formulation, the output projection matrix is removed entirely. The Hadamard matrix is fixed and parameter-free, while the affine rescaling introduces parameters. As a result, the total attention parameter count becomes
( parameters from K,Q,V and after ignoring bias scaling parameters.)
The relative parameter reduction is thus
corresponding to an approximate 25% reduction in attention parameters per multi-head attention block.
@@ 4.3 Computational Complexity
| Property | Dense | Hadamard (FWHT) |
@@
| Forward FLOPs | ||
| Backward FLOPs | ||
@@
| Forward speedup | ||
| Numerical stability | depends on | orthogonal |
@@ 5.1 Experimental Setup
Hardware and Implementation.
All experiments are conducted on 8 × NVIDIA H100 (80GB) GPUs under identical hardware conditions across all model variants to ensure fair comparison. All models are implemented in PyTorch, building upon the NanoGPT codebase Karpathy (2022) (with rope and swiglu) with modifications confined exclusively to the attention module. No additional architectural changes are introduced unless explicitly stated.
Training Configuration
Training is performed in mixed-precision bfloat16 (bf16) using Distributed Data Parallel (DDP) across all 8 GPUs, with global batch sizes reported in tokens. All models are trained with a fixed context length of nctx=1024. We use the AdamW optimizer with b1=0.9, b2=0.95, e=10-8, weight decay of 0.1, and gradient clipping at 1. The learning rate follows a cosine decay schedule with linear warmup; per-model learning rates are reported in Table 2.
| Model Name | Batch Size | Learning Rate | |||||
| Tiny | 124M | 12 | 768 | 12 | 64 | 0.25M | |
| Small | 354M | 24 | 1024 | 16 | 64 | 0.25M | |
| Base | 757M | 24 | 1536 | 16 | 96 | 0.50M |
@@ 5.2 Models Compared
Baseline Models.
The baseline models use the original multi-head attention formulation with a standard output projection layer. We consider 3 baseline configurations corresponding to different model scales. In addition to the size of the model, all architectural components and training settings remain consistent across baseline variants.
Proposed Models.
The proposed models are derived from the NanoGPT baseline by modifying only the attention module. Specifically, we replace the standard dense output projection in the multi-head attention block with the proposed Hadamard-based head mixing mechanism. All other architectural components are kept identical to the baseline. This controlled modification ensures that any observed differences in parameter count, computational cost, memory usage, or performance can be directly attributed to the proposed attention design.
5.3 Efficiency Results
@@
We evaluate our method against the baseline along two primary axes: inference efficiency—throughput, latency, and peak memory—and predictive performance on downstream benchmarks. All efficiency experiments are conducted on a single NVIDIA H100-80GB GPU in BF16 precision unless otherwise stated.
5.3.1 Prefill Phase
@@
Figures 5 and 6 report prefill-phase results. Our method reduces latency and improves throughput across all evaluated model scales and sequence lengths. Speedup is most pronounced at large batch sequence length configurations, where the memory-bandwidth-bound regime dominates and structural parameter reduction yields the greatest benefit.
5.3.2 Decode Phase
Figure 7 presents decode-phase results. Our method achieves consistent throughput improvements that grow with batch size, consistent with the memory-bandwidth-bound nature of autoregressive decoding.
5.3.3 Parameter Reduction
Replacing standard multi-head attention (MHA) with our proposed attention mechanism reduces total parameter count consistently across all three model scales (Small, Medium, and Large), yielding approximately 7% aggregate savings. Notably, the absolute reduction grows with model size, indicating that efficiency gains scale favorably as models increase in depth and width. These results align with the theoretical expectation of removing approximately 25% of attention parameters while leaving feedforward layers unchanged.
5.3.4 GPU Memory Usage
Table 3 summarizes peak GPU memory allocation alongside other decode-phase metrics. Our method reduces peak memory consumption consistently across all model sizes, directly enabling larger batch sizes within the same hardware budget—an important practical benefit for high-throughput serving workloads.
5.3.5 Downstream Evaluation
Our method achieves comparable accuracy to the baseline across several benchmarks, including PIQA Bisk et al. (2019), HellaSwag Zellers et al. (2019), ARC-Easy Clark et al. (2018), and BLiMP Warstadt et al. (2023), with marginal improvements on several tasks.
5.4 Scaling Studies
Validity of Untrained Large-Scale Variants.
The Medium, Large, XL, and XXL configurations were not fully trained due to computational constraints. Crucially, decode-phase efficiency metrics—peak memory, latency, and throughput—are independent of weight values. Under the decode setting, inference is memory-bandwidth bound: performance is governed by parameter count, tensor layout, KV cache footprint, and kernel execution characteristics, not by learned representations. Benchmarking with randomly initialized weights therefore yields valid and reproducible systems-level measurements, consistent with established practice in hardware-oriented decode profiling. The purpose of these larger configurations is to characterize the scaling behavior of structural efficiency, not model quality.
Scaling Behavior.
Throughput gains increase monotonically with model size, from negligible improvement at Tiny scale to at XXL (Table 3). The marginal regression at Tiny is expected: at small scale, kernel launch overhead and fixed runtime constants dominate execution time, masking the memory-bandwidth benefit. As model size grows, the proportion of time spent on memory transfers increases, and our structural reduction yields progressively greater benefit. The same trend holds along the batch and context-length axes: larger batches place more concurrent KV caches in HBM, and longer prompts produce larger per-sequence KV caches, both of which increase the relative share of memory traffic attributable to attention—the component our method directly optimizes.
| Size | Params | Peak Mem (MB) | Latency (ms) | Throughput (tok/s) | ||||||||
| Base | Ours | Base | Ours | Base | Ours | Base | Ours | |||||
| Tiny | 124M | 117M | 5.6% | 3454 | 1.0% | 275 | 1.1% | 59478 | 1.3% | |||
| Small | 354M | 329M | 7.1% | 8525 | 1.1% | 635 | 3.6% | 25797 | 3.8% | |||
| Base | 757M | 701M | 7.4% | 13317 | 1.6% | 1172 | 4.0% | 13974 | 4.1% | |||
| Medium | 1.7B | 1.6B | 5.9% | 24551 | 2.0% | 1982 | 4.5% | 8266 | 4.7% | |||
| Large | 2.6B | 2.4B | 7.7% | 32336 | 2.5% | 2883 | 5.8% | 5683 | 6.2% | |||
| XL | 3.9B | 3.6B | 7.7% | 48030 | 2.5% | 4280 | 5.8% | 3827 | 6.2% | |||
| XXL | 5.6B | 5.1B | 8.9% | 60584 | 2.8% | 5551 | 6.2% | 2951 | 6.6% | |||
| Avg. | — | — | 7.4% | — | — | 2.0% | — | — | 4.3% | — | — | 4.6% |
Naïvely reducing the embedding dimension or FFN expansion ratio produces irregular matrix sizes that misalign with Tensor Core–optimized dimensions (e.g., multiples of 64/128/256), degrading hardware utilization and potentially increasing latency despite fewer parameters. Our method reduces attention parameters while preserving GPU-aligned tensor shapes throughout, ensuring that hardware utilization does not regress. hence we didnt compare our models with vanailla varients
6 Limitations and Discussion
While the proposed Hadamard-based head-mixing layer reduces the theoretical number of operations, there remains significant scope for optimization in its practical implementation. In particular, highly optimized GEMM kernels benefit from decades of engineering effort, whereas our current Hadamard transform and scaling implementation is relatively naive. As a result, despite the reduced arithmetic complexity in theory, our experiments exhibit slightly longer training times compared to the theoritical gain it should have got. We expect that with optimized kernels and better hardware-aware implementations, the practical efficiency of the proposed approach can more closely match its theoretical advantages.
@@ 7 Conclusion
We have proposed an alternative head-mixing mechanism for multi-head attention that replaces the standard dense c×c projection with a fixed Hadamard-based layer. While this modification affects only a single layer, its impact on parameter count, memory usage, and computational cost grows with model width and scale. As models become larger, the cumulative savings become increasingly significant, enabling more efficient inference and facilitating the deployment of attention-based models in performance- and memory-constrained settings.
Acknowledgments
We thank Merai Private Limited for providing GPU resources used in our experiments.

### 30 Mar 2026 v2

@@ Abstract
The dense output projection in multi-head attention scales quadratically with model dimension, contributing significantly to parameter count, memory footprint, and inference cost. We propose replacing this projection with a fixed, parameter-free Walsh–Hadamard Transform (WHT) followed by a diagonal affine transformation. This approach eliminates approximately 25% of attention parameters per block while maintaining global cross-head interaction through an orthogonal, norm-preserving transformation. Our results demonstrate that WHT-augmented models exhibit a steeper validation loss curve relative to training FLOPs compared to dense baselines, suggesting superior compute utilization during training. Crucially, we show that efficiency gains including reduced memory footprint and increased throughput grow monotonically with model size, batch size, and sequence length. We evaluate performance across both prefill and decoding stages, finding that the structured transform consistently outperforms dense projections as complexity increases. Our findings indicate that replacing dense projections with structured transforms allows for more compute-efficient architectures that achieve lower loss than dense models at an equivalent training budget. Our code is available at https://github.com/Shubham2376G/HadamardAttention/tree/main
@@ 3 Motivation
The Hadamard matrix presents a principled answer to the problem of projection overhead. As a fixed, parameter-free orthogonal transform, the Walsh–Hadamard Transform (WHT) uniformly mixes all input dimensions through a butterfly-structured computation (Figure 2), requiring no learned weights while preserving the norm of its input. Crucially, this structured mixing is not merely a computational convenience; it imposes a specific inductive bias on head interaction.
Because the WHT couples every head output with every other through a fixed, maximally spread orthogonal basis, it encourages the model to distribute information across heads in a globally coherent manner. Rather than permitting arbitrary redundancy through unconstrained linear recombination, this structure implicitly regularizes the attention module: heads are incentivized to learn complementary, non-overlapping representations, as only such representations can be efficiently preserved and distinguished under a fixed orthogonal mixing. Our approach aligns with the rationale for orthogonal initialization in deep networks: structured inductive biases that encourage representational diversity in the absence of explicit supervision.
From a computational perspective, the WHT admits an butterfly factorization, replacing the dense matrix multiplication with a cascade of lightweight addition and subtraction operations. Replacing the output projection with this transform eliminates a parameter matrix that typically accounts for approximately 25% of the total attention parameters in a standard Multi-Head Attention (MHA) block, without introducing new hyperparameters or architectural complexity.
In this work, we propose replacing the dense output projection with a structured transformation based on the Hadamard matrix. By doing so, we preserve the essential role of head interaction and dimensional consistency required for residual connections, while significantly reducing the number of learnable parameters. As illustrated in Figure 2, (showing a three-stage butterfly computation for ), this approach challenges the assumption that full dense mixing is necessary for effective head aggregation, suggesting that standard projections are unnecessarily overparameterized.
@@ 4.1 Architecture Overview
We propose a modified Multi-Head Attention (MHA) block in which the standard dense output projection is replaced by a structured, parameter-free mixing operation based on the Walsh-Hadamard Transform (WHT). All other components, including the query (), key (), and value () projections, as well as the per-head attention computation, remain identical to the transformer baseline.
Given an input , the concatenated outputs of the attention heads are denoted as . In a standard transformer, these are mixed via a learned dense projection . We instead formulate the output as:
where is the normalized Walsh-Hadamard matrix satisfying , and denotes the Hadamard (element-wise) product. The vectors represent trainable diagonal scaling and bias parameters, respectively.
@@ 4.2 Parameter Efficiency Analysis
A standard Multi-Head Attention (MHA) block is dominated by four weight matrices: the query (), key (), value (), and output () projections. In a baseline Transformer, each matrix contributes parameters, totaling:
In our proposed architecture, the parameters of the dense matrix are eliminated. The Walsh-Hadamard Transform is parameter-free, requiring only a learned diagonal scaling vector . The resulting parameter count for the attention block becomes:
The relative reduction in attention parameters is calculated as:
Simplifying the expression yields:
For standard model dimensions (e.g., ), the term is negligible. This modification thus provides a constant 25% reduction in attention parameters per block, significantly lowering the model’s memory footprint without reducing the dimensionality of the hidden states.
@@ 4.3 Computational Complexity
| Property | Dense Projection | Hadamard Projection |
@@
| Computational Complexity | ||
| Backward Complexity | ||
@@
| Theoretical Speedup | ||
| Numerical Stability | Data-dependent | Orthogonal |
@@ 5.1 Experimental Setup
Hardware and Implementation. All experiments are conducted on a cluster of NVIDIA H100 (80GB) GPUs. To ensure a fair comparison, all model variants are evaluated under identical hardware conditions. The models are implemented in PyTorch, extending the NanoGPT framework Karpathy (2022) with modern architectural enhancements including Rotary Positional Embeddings (RoPE) and SwiGLU activation functions. Modifications are confined exclusively to the attention module; no additional architectural changes are introduced unless explicitly stated.
Training Configuration. Training is performed in mixed-precision bfloat16 (bf16) using Distributed Data Parallel (DDP). All models are trained with a fixed context length of . We utilize the AdamW optimizer with , , and , applying a weight decay of 0.1 and gradient clipping at 1.0. The learning rate follows a cosine decay schedule with a linear warmup phase. Detailed architectural specifications and training hyperparameters for each model variant are summarized in Table 2.
max width= Model Batch Size Learning Rate Tiny 124M 117M 12 768 12 0.25M Small 354M 329M 24 1024 16 0.25M Base 757M 701M 24 1536 16 0.50M Large 1.3B 1.2B 24 2048 16 1.00M
@@ 5.2 Models Compared
Baseline Configurations
The baselines utilize the standard multi-head attention (MHA) formulation with a dense output projection layer. We evaluate distinct model scales to characterize scaling behavior. All architectural hyperparameters, excluding the projection mechanism, are held constant across variants to isolate the impact of our proposed modification.
WHT-Augmented Models
Our proposed models are identical to the baselines except for a single controlled modification: the dense output projection is replaced by our Hadamard-based head-mixing mechanism. This "drop-in" replacement strategy ensures that observed improvements in parameter efficiency, computational throughput, and scaling performance are directly attributable to the structured orthogonal mixing.
| Model | #Param | Training FLOPs | PPL |
| Small | 334M | 1.55e+19 | 21.54 |
| Ours | 329M | 1.55e+19 | 21.39 |
| Base | 757M | 6.61e+19 | 17.41 |
| Ours | 701M | 6.61e+19 | 17.35 |
| Large | 1.3B | 2.06e+20 | 16.17 |
| Ours | 1.2B | 2.06e+20 | 15.70 |
5.3 Efficiency and Scaling Results
@@
We evaluate the proposed WHT-augmented attention mechanism across two primary axes: (i) Predictive Performance on standard downstream language modeling benchmarks, and (ii) Inference Efficiency, encompassing throughput, latency, and peak memory utilization. All efficiency metrics are measured on a single NVIDIA H100 (80GB) GPU in bfloat16 precision.
5.3.1 Downstream Task Performance
To ensure that the structural constraints of the Hadamard projection do not degrade model expressivity, we evaluate our method on a suite of zero-shot benchmarks: PIQA Bisk et al. (2019), HellaSwag Zellers et al. (2019), ARC-Easy Clark et al. (2018), and BLiMP Warstadt et al. (2023).
As shown in Figure 6, our method achieves comparable accuracy to the baseline across all configurations. This confirms that the orthogonal inductive bias of the Hadamard transform effectively preserves the representational capacity of the attention block while significantly reducing the parameter budget.
5.3.2 Parameter and Memory Efficiency
Parameter Scaling.
As shown in the table 2, the WHT-augmented models yield a consistent reduction in total parameter count, with the relative savings () scaling monotonically with the model dimension. For the Large variant, our method eliminates over 100M parameters, representing a 7.7% aggregate reduction while preserving the core Transformer dimensions.
GPU Memory Utilization.
5.3.3 Inference Throughput and Latency
We stress-test the implementation across the two primary stages of LLM inference: the compute-bound prefill phase and the memory-bandwidth-bound decode phase.
Prefill Phase
@@ 5.3.3 Inference Throughput and Latency
Table 3 and Figures 7 report results for the prefill stage. The reduction in arithmetic complexity from to results in throughput gains of up to 7.0% for the Large model variant. These gains are most pronounced in large-batch and long-sequence configurations, where the memory-bandwidth-bound regime dominates.
max width= Size Model Latency (ms/tok) Throughput (tok/s) Mem (MB) mean std mean std Tiny Baseline 33.157 0.029 150,126 29 16,264 Hadamard 32.366 0.011 156,133 13 16,227 Small Baseline 84.699 0.031 53,154 5 21,330 Hadamard 83.204 0.019 55,956 14 21,233 Base Baseline 129.595 0.074 26,921 13 26,504 Hadamard 126.442 0.052 28,585 15 26,289 Large Baseline 174.158 0.055 16,188 8 32,240 Hadamard 169.501 0.047 17,318 11 31,857
Decode Phase.
For autoregressive decoding (Table 4), our method achieves consistent throughput improvements (up to 3.8%) that scale with batch size. As illustrated in Figure 8, the structured projection improves memory-bandwidth utilization in the single-token generation regime (), which is the primary bottleneck in production environments.
max width= Size Model Latency (ms/tok) Throughput (tok/s) Mem (MB) mean std mean std Tiny Baseline 39.662 0.033 52,820 49 7,299 Hadamard 39.002 0.007 53,730 7 7,259 Small Baseline 105.232 0.218 19,978 1 16,625 Hadamard 102.998 0.229 20,409 12 16,529 Base Baseline 197.373 0.070 10,723 9 25,540 Hadamard 189.879 0.241 11,135 2 25,324 Large Baseline 262.405 0.449 8,025 5 35,016 Hadamard 252.954 0.085 8,312 4 34,632
6 Discussion and Limitations
While the Hadamard-based head-mixing layer significantly reduces the theoretical operation count, its practical latency is currently constrained by implementation overhead. Standard General Matrix Multiply (GEMM) kernels benefit from decades of hardware-level optimization and operator fusion, whereas our current Fast Walsh-Hadamard Transform (FWHT) and affine scaling implementation utilizes relatively unoptimized primitives. Consequently, the empirical training speedups do not yet fully reflect the asymptotic complexity gains. We anticipate that specialized hardware-aware kernels, such as those implemented in Triton or custom CUDA, would allow the practical throughput of our approach to more closely match its theoretical advantages. Furthermore, exploring the interaction between orthogonal mixing and different normalization techniques (e.g., RMSNorm) remains a promising avenue for future work.
@@ 7 Conclusion
In this work, we proposed an efficient head-mixing mechanism for Multi-Head Attention that replaces the dense output projection with a parameter-free Hadamard-based transform followed by a lightweight affine layer. Although this modification is confined to a single component of the attention block, its impact on parameter efficiency and memory footprint scales favorably with model width. As Transformer architectures continue to grow in dimension, the cumulative savings provided by structured projections become increasingly vital. Our results suggest that fully dense mixing is not a prerequisite for effective head aggregation, offering a more sustainable path for deploying high-performance models in resource-constrained environments.